OpenStudy (yuliacosta111):

How many triangles can be made if one angle is 95° and another angle is acute? 1 2 More than 2 None

OpenStudy (yuliacosta111):

@MatthewPFDS

OpenStudy (anonymous):

An acute angle is less than 90

OpenStudy (anonymous):

an obtuse angle is greater than 90 but less than 180. so lets call the obtuse angle 90+x where x lies between 0 and 90. You are told that one of the angles is 95, so the total of these two angles is: 95 + 90 + x = 185 + x what do you know about the sum of the angles in a triangle?

OpenStudy (yuliacosta111):

he sum of measures of three angles of any triangle is equal to the straight angle, also expressed as 180°,

OpenStudy (yuliacosta111):

*the

OpenStudy (anonymous):

The sum of the 3 angles in the triangle will be 180 degrees. An acute angle is from bigger than zero to less than 90 degrees

OpenStudy (anonymous):

Suppose 95 - 30 = 55 95-45 = 50 95 - 5 = 90 it goes on n on

OpenStudy (anonymous):

Infinite.

OpenStudy (anonymous):

That's more than 2

OpenStudy (yuliacosta111):

thxs